# Self-Supervised Skeleton-Based Action Representation Learning: A Benchmark and Beyond

### v2

@@ 1 Introduction
Human activity understanding is an essential topic in the research of computer vision due to its wide applications in real life, such as human-robotics interaction [36], autonomous driving [5], and healthcare [47]. Among the different data modalities for actions, skeletons represent the human body by 3D coordinates of key body joints, which are lightweight, compact, and more robust to changes of view and background. Owing to these desirable advantages, skeleton has attracted much attention in human action analysis.
In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [86, 63, 11, 59]. However, these supervised methods heavily rely on huge amounts of labeled data, which requires time-consuming and expensive data annotation work, limiting the wide applications in the real world. As a remedy to this problem, self-supervised learning (SSL) attracts much attention and has been proven successful for representation learning. It exploits supervisory signals from unlabeled data, learning meaningful prior features and boosting generalization capacity of model for downstream tasks. Motivated by recent success in the image domain, great interest has arisen in adopting SSL for skeleton. However, it is not trivial to transfer these approaches into the skeleton data directly, which are with a more compact spatial structure, additional temporal dimension, and the absence of the background clues. To this end, researchers have made valuable exploration for skeleton-based SSL.
@@
Based on our literature review, it is noticed that most SSL methods for skeleton focus on the single paradigm, learning representations of single granularity, e.g., joint-level features (by masked skeleton modeling, MSM) [100, 80] or sequence-level features (by contrastive learning) [38, 52, 16, 97]. This limits the generalization capacity of the model to more downstream tasks. Although some works [40, 68, 78] make valuable efforts to combine different paradigms, they only achieve mediocre improvement due to the inherent gap of feature modeling mechanisms between the contrastive learning and masked modeling [57, 23]. To this end, a novel SSL approach for skeleton is proposed to fully boost the generalization capacity of SSL model, which integrates the contrastive learning and MSM to learn the joint, clip, and sequence level representations jointly. Specifically, we fully utilize the novel motion pattern exposed by the manual designed augmentations and model training for sequence-level contrastive representation learning, while adopting MSM for the joint-level feature modeling. Besides, we further propose a novel clip-level contrastive learning method, which significantly boosts the short-term model capacity, along with an effective post-distillation strategy to achieve a more compact representation space. Finally, extensive experiments under five downstream tasks, not limited to the single action recognition task used in most previous works, demonstrate the promising generalization capacity of the proposed method.
@@
  • We present a comprehensive benchmark of existing skeleton SSL works as well as a summary of the popular datasets and downstream tasks as well as the insightful analysis from the perspective of model backbones, pre-training paradigms, and future possible directions. Finally, we demonstrate promising performance of our proposed method on five downstream tasks for versatile action representation learning.
The remaining sections are organized as follows: We first present a thorough review in Sec. 2, for skeleton-based SSL representation learning. Subsequently, based on our investigation, we propose a new method exploring the combination of contrastive learning and masked modeling tasks in Sec. 3. Then, we comprehensively benchmark existing methods in Sec. 4, and verify the effectiveness of our proposed approach. Finally, we conclude and summarize with possible future directions in Sec. 5.
@@ 2 Review on Skeleton-Based Action Representation SSL
Generally, a two-stage paradigm is utilized in skeleton-based SSL, i.e., pre-training on pretext tasks first and then fine-tuning on downstream tasks. In the pre-training stage, different pretext tasks are designed for deep neural networks, capturing training signals derived from the data itself, called the process of self-supervision [66]. After that, the learned knowledge as feature representations is transferred to downstream tasks as shown in Fig. 1 (c). In principle, this part is not only the goal of SSL, i.e., to improve downstream task performance with learned representations, but also the way to assess the quality of representation learning methods. Note that some downstream tasks, e.g., motion prediction and 2D-to-3D lifting, are not considered in this survey, because they essentially do not rely on the supervisory signal of human annotation and can serve as pre-training pretext tasks themselves, leading to possible unfair comparison. For reviews on these topics, we direct readers to [50, 45].
@@ 2.1 Human Skeleton Representation
Skeleton represents a human motion as a sequence of the body keypoint coordinates over time. The collection methods of skeleton can be divided into two main categories, i.e., marker-based and markerless methods, shown in Fig. 1 (a).
Marker-based motion capture (Mocap) systems often rely on inertial measurement units (IMU) or reflective markers, placed on the body to track the movement of humans. It can provide reliable skeleton data with even sub-millimeter accuracy [4]. However, this method is too costly in many application scenarios and requires highly trained personnel to operate. Meanwhile, it suffers from great inconvenience, e.g., the time-consuming placement process and a requirement for a controlled environment. The datasets [31, 75] for generation tasks, e.g., pose estimation, are usually collected in this way to guarantee the accuracy.
Markerless Mocap often depends on deep learning algorithms for pose estimation from RGB and depth data. Video RGB data can be easily obtained from the Internet while we can also utilize the camera hardware, e.g., the depth camera including Microsoft Azure Kinect, and single or multiple RGB video cameras, to collect the RGB, infrared, and depth images in deployment. Then to get the 3D motion, computer vision algorithms, on multi-view geometry and pose estimation, i.e., OpenPose [6] are then employed to detect and extract joint center locations. The whole pipeline is presented in Fig. 1 (a). However, due to the limitations of hardware and estimation algorithms, such method can have large errors compared to marker-based methods [15]. Nevertheless, it is still chosen for most skeleton-based datasets [62, 44, 43] for its simpleness and convenience.
@@ 2.2.1 Context-Based Methods
Context-based methods generate the supervisory training signals according to the inherent contextual information of provided data. The model is encouraged to learn the spatial-temporal relationships by training on the pre-defined task. Emerging from the image domain, the pretext tasks rely on the context understanding, e.g., rotation prediction [21]. In contrast, skeleton data introduces an additional temporal dimension, and possesses more compact spatial information, which presents a new challenge on how to mine the meaningful spatial-temporal context of skeleton by pretext tasks. Typically, there are three common types of context-based pretext tasks, view-invariance-based, temporal-order-based, and motion-prior-based method as shown in Fig. 3.
@@
1) View Invariance. Due to the variance of the observation viewpoint, the skeleton estimation can suffer from occlusion and noise. Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [19]. Li et al. [37] proposed a view classification pretext task for unsupervised action representation learning. Specifically, as shown in Fig. 3 (a), skeleton sequences of different views are fed into the encoder, subsequent to which a view classifier predicts the view labels. To learn the view-invariant features, a Gradient Reversal Layer [18] is added to reverse the optimization direction of the encoder. Resort to this, the encoder can learn the features insensitive to view in an adversarial manner. Likewise, Paoletti et al. [55] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
There are also some subsequent SSL works [87, 53, 3] focusing on learning view-invariant representations. However, they adopted a contrastive learning paradigm to explicitly learn the alignment between the two views and showed a significant performance improvement.
@@
Some skeleton-based SSL works [40, 83] are equipped with temporal sorting to model the temporal dependencies, similar to SSL works in video [82, 34]. However, such pretext task is often integrated with other pretext tasks, because it only models temporal features explicitly, leaving the crucial spatial representations of skeleton under-explored. Therefore, new pretext tasks are studied to jointly learn the spatial-temporal relationships of skeleton.
3) Motion Prior. The motion dynamics of skeleton joints contain rich semantic information beneficial for action understanding. Therefore, researchers propose to leverage the inherent motion prior knowledge to generate pseudo labels for pre-training as shown in Fig. 3 (c). Cheng et al. [12] designed a movement direction prediction task. The model is constrained to estimate the direction of the instantaneous joint velocity, i.e., whether the targeted joints are moving in a positive direction. Based on this, Kim et al. [33] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning. The key idea is to predict the motion direction and magnitude of the central joint and other joints during a random interval. On the other hand, Yang et al. [91] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation. The model is pre-trained by a regression task to predict this prior knowledge.
@@ 2.2.2 Generative Learning Methods
When $\mathcal{T}_{in}$ and $\mathcal{T}_{tar}$ are both the identity function, the model is constrained to reconstruct the input data simply, known as AutoEncoder (AE). It naturally creates an information bottleneck to achieve dimension reduction, mapping data from input space onto low-dimensional feature space. Due to its simpleness, AE is used to learn prior representations in earlier skeleton-based works [68, 84]. However, since the skeleton data is still redundant in spatial and temporal dimensions, the encoder tends to memorize the input data at a low level instead of modeling the high-level semantic knowledge. Then, denoising AE (DAE) is studied for alleviating this shortcut [69, 54]. For example, Nie et al. [54] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features. On the other hand, inspired by the remarkable success of masked language/image modeling (MLM, MIM), which can be regarded as a variant of DAE, researchers have aroused a surge of interest in exploring masked skeleton modeling (MSM) for human action representation learning. The earlier works are mainly based on recurrent neural network (RNN) or graph convolutional network (GCN) to predict the masked skeletons. However, lacking good scalability, these models only show mediocre performance improvement. Recently, inspired by the success of Vision Transformer (ViT) as masked autoencoders (MAE) [25], Transformer-based models have been explored for the masked skeleton modeling [9, 51]. However, due to the relative redundancy in skeleton and the lack of a large-scale dataset as ImageNet, directly applying Transformer for masked skeleton reconstruction can suffer from the over-fitting problem. To this end, researchers have made endeavors on two crucial designs, masking strategy and reconstruction target, which are discussed as follows.
@@
1) Masking Strategy has been proved crucial as the design of $\mathcal{T}_{in}$ in Eq. (1) for MIM. In the field of RGB images, MAE [25] adopts a patch-based masking strategy with a large masking ratio of 75%. For the skeleton data, researchers have fully explored the masking strategy in the spatial-temporal dimension. Although the optimal masking strategy and ratio can differ with the encoder backbone and MSM task setting, most works have found that spatially body-part-based and temporally segment-based strategy with large masking ratio, e.g., 90% in [52], can produce decent results. Concretely, body-part level masking refers to performing masking regarding the different joints in a body part as a whole, e.g., torso and left leg, while the temporal segment-based strategy masks the same joints across consecutive frames. These designs aim to reduce the shortcuts caused by spatial-temporal redundancy in skeleton, which is in line with the observation in images [26] and videos [74].
@@
Recently, more elaborate masking strategies have been studied to achieve more effective exploitation of valuable semantic information. The work [85] found that masking the limbs is always better than masking the torso and head, especially the right hand and leg. MAMP [51] proposes a motion-aware masking strategy. The moving parts are located and masked by calculating the motion displacement between adjacent frames, which achieves better results than the random masking strategy. These results demonstrate that masking motion regions, which are often semantic-rich, promotes the model to learn more meaningful features in masked motion modeling. It can be also explained by the theory in [99], i.e., because samples from the same action often contain similar motion patterns, this masking strategy on the motion regions implicitly leads to better alignment of mask-induced positive pairs, achieving more discriminative feature spaces.
@@
Meanwhile, some literature adopts special mask designs, yielding new MSM task forms. MS2[40] directly utilizes temporal frame-level masks to train the model on the motion prediction task. MotionBERT [103] adopts a 2D-to-3D lifting pretext task, in which it masks all the values in depth $z$ channel and partially in $x$ and $y$ channels, encouraging the model to predict the original 3D skeletons. These works can be divided into masked skeleton modeling in a broad sense.
@@
2) Reconstruction Target, known as the implementation of $\mathcal{T}_{tar}$, also varies from the input space to the feature space. Most existing works reconstruct the skeleton data in the input coordinate space using MSE loss. Some works [78] propose to perform temporally reverse reconstruction to learn the skeleton dynamics instead of trivial representations that just remember the input.
@@
Instead of directly reconstructing the input, Mao et al. [51] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames. The work [90] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space. This mapping function is artificially defined according to the spatial-temporal relationship of skeleton joints. Therefore, the model can learn the spatial relation and temporal dependency by reconstructing the color of the skeleton cloud.

### v3

@@ 1 Introduction
Human activity understanding is an essential topic in the research of computer vision due to its wide applications in real life, such as human-robotics interaction [lee2020real], autonomous driving [camara2020pedestrian], and healthcare [lopez2019human]. Among the different data modalities for actions, skeletons represent the human body by 3D coordinates of key body joints, which are lightweight, compact, and more robust to changes of view and background. Owing to these desirable advantages, skeleton has attracted much attention in human action analysis.
In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [yan2018spatial, shi2019two, cheng2020skeleton, rao2024hierarchical]. However, these supervised methods heavily rely on huge amounts of labeled data, which requires time-consuming and expensive data annotation work, limiting the wide applications in the real world. As a remedy to this problem, self-supervised learning (SSL) attracts much attention and has been proven successful for representation learning. It exploits supervisory signals from unlabeled data, learning meaningful prior features and boosting generalization capacity of model for downstream tasks. Motivated by recent success in the image domain, great interest has arisen in adopting SSL for skeleton. However, it is not trivial to transfer these approaches into the skeleton data directly, which are with a more compact spatial structure, additional temporal dimension, and the absence of the background clues. To this end, researchers have made valuable exploration for skeleton-based SSL.
@@
Based on our literature review, it is noticed that most SSL methods for skeleton focus on the single paradigm, learning representations of single granularity, e.g., joint-level features (by masked skeleton modeling, MSM) [zheng2018unsupervised, wu2022skeletonmae] or sequence-level features (by contrastive learning) [li20213d, mao2022cmd, hico2023, zhang2022hiclr]. This limits the generalization capacity of the model to more downstream tasks. Although some works [lin2020ms2l, su2020predict, wang2022contrast] make valuable efforts to combine different paradigms, they only achieve mediocre improvement due to the inherent gap of feature modeling mechanisms between the contrastive learning and masked modeling [qi2023contrast, gui2023survey]. To this end, a novel SSL approach for skeleton is proposed to fully boost the generalization capacity of SSL model, which integrates the contrastive learning and MSM to learn the joint, clip, and sequence level representations jointly. Specifically, we fully utilize the novel motion pattern exposed by the manual designed augmentations and model training for sequence-level contrastive representation learning, while adopting MSM for the joint-level feature modeling. Besides, we further propose a novel clip-level contrastive learning method, which significantly boosts the short-term modeling capacity, along with an effective post-distillation strategy to achieve a more compact representation space. Finally, extensive experiments under five downstream tasks, not limited to the single action recognition task used in most previous works, demonstrate the promising generalization capacity of the proposed method.
@@
  • We present the first multi-task benchmark of existing skeleton SSL works with the insightful analysis from the perspective of model backbones, pre-training paradigms, and future possible directions. Finally, we demonstrate promising performance of our proposed method on five downstream tasks for versatile action representation learning.
The remaining sections are organized as follows: We first present a thorough review in Sec. 2, for skeleton-based SSL representation learning. Subsequently, based on our investigation, we propose a new method exploring the combination of contrastive learning and masked modeling tasks in Sec. 3. Then, we comprehensively benchmark existing methods in Sec. 4 on multiple downstream tasks, and verify the effectiveness of our proposed approach. Finally, we conclude and summarize with possible future directions in Sec. 5.
@@ 2 Review on Skeleton-Based Action Representation SSL
Generally, a two-stage paradigm is utilized in skeleton-based SSL, i.e., pre-training on pretext tasks first and then fine-tuning on downstream tasks. In the pre-training stage, different pretext tasks are designed for deep neural networks, capturing training signals derived from the data itself, called the process of self-supervision [shurrab2022self]. After that, the learned knowledge as feature representations is transferred to downstream tasks as shown in Fig. 1 (c). In principle, this part is not only the goal of SSL, i.e., to improve downstream task performance with learned representations, but also the way to assess the quality of representation learning methods. Note that some downstream tasks, e.g., motion prediction and 2D-to-3D lifting, are not considered in this survey, because they essentially do not rely on the supervisory signal of human annotation and can serve as pre-training pretext tasks themselves, leading to possible unfair comparison. For reviews on these topics, we direct readers to [lyu20223d, liu2022recent].
@@ 2.1 Human Skeleton Representation
The collection methods of skeleton can be divided into two main categories, i.e., marker-based and markerless methods, shown in Fig. 1 (a).
Marker-based motion capture (Mocap) systems often rely on inertial measurement units (IMU) or reflective markers, placed on the body to track the movement of humans. It can provide reliable skeleton data with even sub-millimeter accuracy [buckley2019role]. However, this method is too costly in many application scenarios and requires highly trained personnel to operate. Meanwhile, it suffers from great inconvenience, e.g., the time-consuming placement process and a requirement for a controlled environment. The datasets [ionescu2013human3, von2018recovering] for generation tasks, e.g., pose estimation, are usually collected in this way to guarantee the accuracy.
Markerless Mocap often depends on deep learning algorithms for pose estimation from RGB and depth data. Video RGB data can be easily obtained from the Internet while we can also utilize the camera hardware, e.g., the depth camera including Microsoft Azure Kinect, and single or multiple RGB video cameras, to collect the RGB, infrared, and depth images in deployment. Then to get the 3D motion, computer vision algorithms, on multi-view geometry and pose estimation, i.e., OpenPose [cao2017realtime] are then employed to detect and extract joint center locations. The whole pipeline is presented in Fig. 1 (a). However, due to the limitations of hardware and estimation algorithms, such method can have large errors compared to marker-based methods [dolatabadi2016concurrent]. Nevertheless, it is still chosen for most skeleton-based datasets [shahroudy2016ntu, liu2019ntu, liu2020pku] for its simpleness and convenience.
@@ 2.2 SSL Methods for Skeleton
SSL has been widely explored in the previous literature for the image domain. Some early works adopt the context-driven pretext tasks to model the inherent data structure, e.g., predicting the rotation label [gidaris2018unsupervised], correlation map [li2023correlational], and clusters based on the pseudo-labels [caron2018deep, zhan2020online]. Besides, contrastive-style methods push the performance of SSL to new heights, including invariance decorrelation [zbontar2021barlow, bardes2021vicreg], distillation-based [grill2020bootstrap, chen2021exploring, caron2021emerging, oquab2023dinov2, darcet2023vision], clustering-based [caron2020unsupervised, li2020prototypical]. Meanwhile, the generative learning, especially the recent masked image modeling [xie2022simmim, he2022masked], has drawn a surge of interest and been proven effective for downstream vision downstream tasks. These works have played an important role and motivate the new efforts for skeleton-based SSL. However, due to the huge difference between human skeleton and image data as discussed in Sec. 1, it is suboptimal to transfer these methods directly for skeleton representation learning. Therefore, we present this literature review with a high-level methodological taxonomy, and also shed light on the challenges and skeleton-specific designs to motivate further work.
@@ 2.2.1 Context-Based Methods
Context-based methods generate the supervisory training signals according to the inherent contextual information of provided data. The model is encouraged to learn the spatial-temporal relationships by training on the pre-defined task. Emerging from the image domain, the pretext tasks rely on the context understanding, e.g., rotation prediction [gidaris2018unsupervised]. In contrast, skeleton data introduces an additional temporal dimension, and possesses more compact spatial information, which presents a new challenge on how to mine the meaningful spatial-temporal context of skeleton by pretext tasks. Typically, there are three common types of context-based pretext tasks, view-invariance-based, temporal-order-based, and motion-prior-based method as shown in Fig. 3.
@@
1) View Invariance. Due to the variance of the observation viewpoint, the skeleton estimation can suffer from occlusion and noise. Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [gao2022global]. Li et al. [li2018unsupervised] proposed a view classification pretext task for unsupervised action representation learning. Specifically, as shown in Fig. 3 (a), skeleton sequences of different views are fed into the encoder, subsequent to which a view classifier predicts the view labels. To learn the view-invariant features, a Gradient Reversal Layer [ganin2015unsupervised] is added to reverse the optimization direction of the encoder. Resort to this, the encoder can learn the features insensitive to view in an adversarial manner. Likewise, Paoletti et al. [paoletti2022unsupervised] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
There are also some subsequent SSL works [yang2022via, men2023focalized, bian2022view] focusing on learning view-invariant representations. However, they adopted a contrastive learning paradigm to explicitly learn the alignment between the two views and showed a significant performance improvement.
@@
Some skeleton-based SSL works [lin2020ms2l, xu2021prototypical] are equipped with temporal sorting to model the temporal dependencies, similar to SSL works in video [xu2019self, lee2017unsupervised]. However, such pretext task is often integrated with other pretext tasks, because it only models temporal features explicitly, leaving the crucial spatial representations of skeleton under-explored. Therefore, new pretext tasks are studied to jointly learn the spatial-temporal relationships of skeleton.
3) Motion Prior. The motion dynamics of skeleton joints contain rich semantic information beneficial for action understanding. Therefore, researchers propose to leverage the inherent motion prior knowledge to generate pseudo labels for pre-training as shown in Fig. 3 (c). Cheng et al. [cheng2021hierarchical] designed a movement direction prediction task. The model is constrained to estimate the direction of the instantaneous joint velocity, i.e., whether the targeted joints are moving in a positive direction. Based on this, Kim et al. [kim2022global] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning. The key idea is to predict the motion direction and magnitude of the central joint and other joints during a random interval. On the other hand, Yang et al. [yang2022motion] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation. The model is pre-trained by a regression task to predict this prior knowledge.
@@ 2.2.2 Generative Learning Methods
When $\mathcal{T}_{in}$ and $\mathcal{T}_{tar}$ are both the identity function, the model is constrained to reconstruct the input data simply, known as AutoEncoder (AE). It naturally creates an information bottleneck to achieve dimension reduction, mapping data from input space onto low-dimensional feature space. Due to its simpleness, AE is used to learn prior representations in earlier skeleton-based works [su2020predict, xu2021unsupervised]. However, since the skeleton data is still redundant in spatial and temporal dimensions, the encoder tends to memorize the input data at a low level instead of modeling the high-level semantic knowledge. Then, denoising AE (DAE) is studied for alleviating this shortcut [su2021self, nie2020unsupervised]. For example, Nie et al. [nie2020unsupervised] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features. On the other hand, inspired by the remarkable success of masked language/image modeling (MLM, MIM), which can be regarded as a variant of DAE, researchers have aroused a surge of interest in exploring masked skeleton modeling (MSM) for human action representation learning. The earlier works are mainly based on recurrent neural network (RNN) or graph convolutional network (GCN) to predict the masked skeletons. However, lacking good scalability, these models only show mediocre performance improvement. Recently, inspired by the success of Vision Transformer (ViT) as masked autoencoders (MAE) [he2022masked], Transformer-based models have been explored for the masked skeleton modeling [chen2022hierarchically, mao2023masked]. However, due to the relative redundancy in skeleton and the lack of a large-scale dataset as ImageNet, directly applying Transformer for masked skeleton reconstruction can suffer from the over-fitting problem. To this end, researchers have made endeavors on two crucial designs, masking strategy and reconstruction target, which are discussed as follows.
@@
1) Masking Strategy has been proved crucial as the design of $\mathcal{T}_{in}$ in Eq. (1) for MIM. In the field of RGB images, MAE [he2022masked] adopts a patch-based masking strategy with a large masking ratio of 75%. For the skeleton data, researchers have fully explored the masking strategy in the spatial-temporal dimension. Although the optimal masking strategy and ratio can differ with the encoder backbone and MSM task setting, most works have found that spatially body-part-based and temporally segment-based strategy with large masking ratio, e.g., 90% in [mao2022cmd], can produce decent results. Concretely, body-part level masking refers to performing masking regarding the different joints in a body part as a whole, e.g., torso and left leg, while the temporal segment-based strategy masks the same joints across consecutive frames. These designs aim to reduce the shortcuts caused by spatial-temporal redundancy in skeleton, which is in line with the observation in images [he2022masked] and videos [tong2022videomae].
@@
Recently, more elaborate masking strategies have been studied to achieve more effective exploitation of valuable semantic information. The work [yan2023skeletonmae] found that masking the limbs is always better than masking the torso and head, especially the right hand and leg. MAMP [mao2023masked] proposes a motion-aware masking strategy. The moving parts are located and masked by calculating the motion displacement between adjacent frames, which achieves better results than the random masking strategy. These results demonstrate that masking motion regions, which are often semantic-rich, promotes the model to learn more meaningful features in masked motion modeling. It can be also explained by the theory in [zhang2022mask], i.e., because samples from the same action often contain similar motion patterns, this masking strategy on the motion regions implicitly leads to better alignment of mask-induced positive pairs, achieving more discriminative feature spaces.
@@
Meanwhile, some literature adopts special mask designs, yielding new MSM task forms. MS2[lin2020ms2l] directly utilizes temporal frame-level masks to train the model on the motion prediction task. MotionBERT [zhu2023motionbert] adopts a 2D-to-3D lifting pretext task, in which it masks all the values in depth $z$ channel and partially in $x$ and $y$ channels, encouraging the model to predict the original 3D skeletons. These works can be divided into masked skeleton modeling in a broad sense.
@@
2) Reconstruction Target, known as the $\mathcal{T}_{tar}$, also varies from the input space [he2022masked, xie2022simmim] to the feature space [assran2023self, gao2024mimic, xie2022masked, haghighat2023pre, bardes2024revisiting] as the image domain. Most existing works reconstruct the skeleton data in the input coordinate space using MSE loss. Some works [wang2022contrast] propose to perform temporally reverse reconstruction to learn the skeleton dynamics instead of trivial representations that just remember the input.
@@
Instead of directly reconstructing the input, Mao et al. [mao2023masked] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames. The work [yang2023self] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space. This mapping function is artificially defined according to the spatial-temporal relationship of skeleton joints. Therefore, the model can learn the spatial relation and temporal dependency by reconstructing the color of the skeleton cloud.